The following data were obtained when an extruded breakfast cereal was analyzed for total fiber

  • Sample wt (mg) 1002.8
  • Residue wt (mg) 151.9
  • Protein wt (mg) 13.1
  • Ash wt (mg) 21.1
  • Blank wt (mg) 6.1
  • Resistant starch (mg) 35.9

What is percent total fiber (a) without and (b) with correction for resistant starch?
Answer #1

Answer:-

According to the above data:-

  • Sample weight (mg) = 1002.8
  • Residue weight (mg) = 151.9
  • Protein weight (mg) = 13.1
  • Ash weight (mg) = 21.1
  • Blank weight (mg) = 6.1
  • Resistant starch (mg) = 35.9

The above data were obtained when an extruded breakfast cereal was analysed for total fiber by AOAC method in which total dietary fibre can be estimated by:- (weight of residue - weight of protein - weight of ash - weight of blank/ sample weight)

(a) The percentage of total fibre without correction for resistant starch

= (151.9 - 13.1 - 21.1 - 6.1/ 1002.8) × 100 = 11.1%

(b) The percentage of total fibre with correction for resistant starch

= (151.9 - 13.1 - 21.1 - 6.1 - 35.9/ 1002.8) × 100 = 7.54%

Hope the above answer comes up to your expectation. Please do come up with your feedback.